# Importance sampling for Sobol’ indices estimation

### v1

@@ 3.2 Optimal asymptotic variance expression using importance sampling
We now investigate the case where the reference distribution of the input vector $\mathbf{X}$ can be freely modified, while the distribution of the noise variable $W$ remains fixed. We define $\phi(\mathbf{X})=\sqrt{\mathbb{E}[Y^{2}|\mathbf{X}]}$, noting that $\phi$ is model-specific and independent of the sampling distribution; its dependence on $p$ or $q$ does not need to be made explicit in the notation. Under $p$, the optimal asymptotic variance for estimating $\eta_{u}$ is given by
$\sigma^{2}_{\mathrm{opt},u}(p)=4\,\mathbb{E}_{p}\left[m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]\right]-3\,\mathbb{E}_{p}\left[m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}.$ (10)
We now derive the expression of the asymptotic variance under an arbitrary auxiliary distribution $q$, and compare it to the reference case.
proposition
Let $q(\mathbf{x})$ be a sampling density over $\mathbf{X}$, and let $Z_{u}$ be defined as in Lemma 1. Then the optimal asymptotic variance is given by
$\sigma^{2}_{\mathrm{opt},u}(q)=4\,\mathbb{E}_{q}\left[\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}\cdot\mathbb{E}_{q}[Z_{u}^{2}\mid\mathbf{X}_{u}]\right]-3\,\mathbb{E}_{q}\left[\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{4}\right]-\eta_{u}^{2}.$ (11)
Using Lemma 1 and substituting the expressions of $Z_{u}$, this variance can be rewritten under $p$ as
$\sigma^{2}_{\mathrm{opt},u}(q)$ $=4\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\cdot m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]\right]$ (12)
$\quad-3\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\cdot m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}.$
This expression highlights how the asymptotic variance depends explicitly on the choice of the sampling distribution $q$ through the weights $w_{u}$ and $w_{\bar{u}}$, and through the structure of $Z_{u}$. Optimizing $q$ can therefore lead to substantial variance reduction.
@@ 3.3 Optimal distribution: sequential construction
We aim to minimize the asymptotic variance $\sigma^{2}_{\mathrm{opt},u}(q)$ by optimizing the components of the sampling distribution $q$ in a sequential manner. Our strategy proceeds in two steps. First, for a fixed marginal $q_{u}$, we determine the optimal conditional density $q^{*}_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})$ that minimizes the variance contribution given $\mathbf{x}_{u}$. This optimal conditional density is then substituted into the variance expression. In the second step, the resulting objective is minimized with respect to the marginal $q_{u}$. To characterize the optimal distribution, we rely on a key variational lemma recalled below.
Lemma
@@
Let $\mathcal{D}_{k}$ denote the set of probability density functions on $\mathbb{R}^{k}$.
Then, for any $g\in\mathcal{D}_{k}$, the value $q\overset{ae}{=}g$ uniquely minimizes the functional
$q\mapsto\int_{\mathbb{R}^{k}}\frac{g^{2}(\mathbf{x})}{q(\mathbf{x})}\,d\mathbf{x}\quad,\quad q\in\mathcal{D}_{k}.$
Proof
Let $q\in\mathcal{D}_{k}$. By Jensen’s inequality applied to the convex function $t\mapsto t^{2}$, we obtain:
[…]
Where equality is achieved if and only if $q=g$ almost everywhere.
First step: optimization of $q_{\bar{u}\mid u}$
We begin by optimizing the conditional component $q_{\bar{u}\mid u}$ of $q$, assuming that $q_{u}$ is fixed. From Proposition 3.2, the conditional distribution influences the asymptotic variance $\sigma^{2}_{\mathrm{opt},u}(q)$ in eq(3.6) only through the term
$\mathbb{E}_{\begin{subarray}{c}p\end{subarray}}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right].$
@@
Minimizing this expectation is therefore sufficient, since it is the only component of the variance that depends on $q_{\bar{u}\mid u}$. Applying Lemma 2 conditionally on $\mathbf{X}_{u}$ gives the optimal conditional density:
$q^{*}_{\bar{u}\mid u}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})=\frac{p_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi(\mathbf{X})}{\mathbb{E}_{\begin{subarray}{c}p\end{subarray}}\left[\phi(\mathbf{X})\mid\mathbf{X}_{u}\right]},$ (13)
@@
where the denominator ensures normalization. With this choice, the variance contribution when $q_{\bar{u}\mid u}=q^{*}_{\bar{u}\mid u}$ becomes $\left(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\right)^{2}$, instead of $\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}]$ when $\mathbf{X}_{\bar{u}\mid u}\sim p_{\bar{u}}$ (i.e., $w_{\bar{u}\mid u}=1$). A full derivation is provided in the Appendix. Since the function $x\mapsto x^{2}$ is convex, Jensen’s inequality confirms that
$\left(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\right)^{2}\leq\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}],$
which guarantees that the optimal conditional distribution $q^{*}_{\bar{u}\mid u}$ systematically reduces this part of the variance.
@@
We now fix the component $q_{\bar{u}\mid u}$, and seek to minimize the asymptotic variance over the marginal $q_{u}$. For any such fixed conditional, the sampling density is of the form $q=q_{u}\cdot q_{\bar{u}\mid u}$, and the corresponding variance reads
$\sigma_{\operatorname{opt},u}^{2}(q)=\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\,S(\mathbf{X}_{u})\right]-\eta_{u}^{2},$
where the integrand $S(\mathbf{x}_{u})$ depends on the choice of $q_{\bar{u}\mid u}$. The optimal marginal minimizing this expression is given by
$q_{u}^{*}(\mathbf{x}_{u})=\frac{p(\mathbf{x}_{u})\,\sqrt{S_{q_{\bar{u}\mid u}}(\mathbf{x}_{u})}}{\mathbb{E}_{p}\left[\sqrt{S_{q_{\bar{u}\mid u}}(\mathbf{X}_{u})}\right]},\qquad\text{and}\qquad\sigma^{2}_{\operatorname{opt},u}(q)=\left(\mathbb{E}_{p}\left[\sqrt{S_{q_{\bar{u}\mid u}}(\mathbf{X}_{u})}\right]\right)^{2}-\eta_{u}^{2}.$ (14)
@@
  • Case A – Reference conditional: $q_{\bar{u}\mid u}=p_{\bar{u}\mid u}$
    $S_{A}(\mathbf{x}_{u})=4\,m_{u}^{2}(\mathbf{x}_{u})\,\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}]-3\,m_{u}^{4}(\mathbf{x}_{u}).$
  • Case B – Optimal conditional: $q_{\bar{u}\mid u}=q^{*}_{\bar{u}\mid u}$
    $S_{B}(\mathbf{x}_{u})=4\,m_{u}^{2}(\mathbf{x}_{u})\,\left(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\right)^{2}-3\,m_{u}^{4}(\mathbf{x}_{u}).$
@@
This satisfies the pointwise inequality $S_{B}(\mathbf{x}_{u})\leq S_{A}(\mathbf{x}_{u})$, shown in step one. The full chain of inequalities then reads
$\sigma^{2}_{\operatorname{opt},u}(q^{*})\leq\sigma^{2}_{\operatorname{opt},u}(q_{u}^{*}\cdot p_{\bar{u}\mid u})\leq\sigma^{2}_{\operatorname{opt},u}(p),$
this follows from Jensen’s inequality:
$\left(\mathbb{E}_{p}[\sqrt{S_{B}(\mathbf{X}_{u})}]\right)^{2}\leq\left(\mathbb{E}_{p}[\sqrt{S_{A}(\mathbf{X}_{u})}]\right)^{2}\leq\mathbb{E}_{p}[S_{A}(\mathbf{X}_{u})].$
@@
That is, joint optimisation in Case B achieves the greatest variance reduction, demonstrating the improved efficiency obtained when both components of $q$ are optimised together.
3.4 Zero-variance optimal distribution
@@
In the idealised case where the function $f$ is fully deterministic with all inputs controllable (i.e., there is no external noise variable $W$ and $Y=f(\mathbf{X})$), it becomes theoretically possible to achieve zero-variance in the estimation of $\eta_{u}$. This setting can be interpreted as a special case of the general framework discussed earlier. We still proceed in two steps, but expressions simplify significantly. In particular, the asymptotic variance now takes the form
$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\left(4\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,Y^{2}\mid\mathbf{X}_{u}\right]-3\,m_{u}^{4}(\mathbf{X}_{u})\right)\right]-\eta_{u}^{2}.$ (15)
Applying Lemma 2 sequentially yields:
$q_{\bar{u}}^{*}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})=\frac{p_{\bar{u}}(\mathbf{x}_{\bar{u}})\,f(\mathbf{x})}{m_{u}(\mathbf{x}_{u})},\qquad q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,m_{u}^{2}(\mathbf{x}_{u})}{\eta_{u}}.$ (16)
@@
$q^{*}(\mathbf{x})=q_{u}^{*}(\mathbf{x}_{u})\cdot q_{\bar{u}}^{*}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})=\frac{p(\mathbf{x})\,f(\mathbf{x})\,m_{u}(\mathbf{x}_{u})}{\eta_{u}},\qquad\text{and}\qquad\sigma^{2}_{\mathrm{opt},u}(q^{*})=0.$ (17)
Hence, the density $q^{*}(\mathbf{x})$ achieves a zero-variance estimator in theory. However, this ideal result is not achievable in practice, as the optimal distribution depends on the unknown quantity $\eta_{u}$ that we aim to estimate.

### v2

@@ 3.2 Optimal asymptotic variance expression using importance sampling
We define $\phi(\mathbf{x})=\sqrt{\mathbb{E}[Y^{2}\mid\mathbf{X}=\mathbf{x}]}$, noting that $\phi$ is model-specific and does not depend on the choice of sampling distribution. Under $p$, the optimal asymptotic variance for estimating $\eta_{u}$ is given by
$\sigma^{2}_{\mathrm{opt},u}(p)=\mathbb{E}_{p}\Big[4\,m^{2}(\mathbf{X}_{u})\,\nu_{p}(\mathbf{X}_{u})+m^{4}(\mathbf{X}_{u})\Big]-\eta_{u}^{2},$ (10)
where the conditional variance $\nu_{p}$ is given by
$\nu_{p}(\mathbf{X}_{u}):=\mathrm{Var}_{p}(Y\mid\mathbf{X}_{u})=\mathbb{E}_{p}[Y^{2}\mid\mathbf{X}_{u}]-m^{2}(\mathbf{X}_{u})=\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}]-m^{2}(\mathbf{X}_{u}),$
We next extend this result to an auxiliary distribution $q$ and compare it with the reference case.
Proposition 3.2
Let $Z_{u}$ be defined as in Lemma 3.1. Then, for the sampling distribution $q$, the optimal asymptotic variance is given by
$\boxed{\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{q}\Big[4\,\big(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\big)^{2}\,\mathrm{Var}_{q}(Z_{u}\mid\mathbf{X}_{u})+\big(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\big)^{4}\Big]-\eta_{u}^{2}.}$ (11)
Substituting the expressions of $Z_{u}$, $\sigma^{2}_{\mathrm{opt},u}(q)$ can be equivalently expressed under $p$ as
$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\Big[w_{u}(\mathbf{X}_{u})\big(4\,m^{2}(\mathbf{X}_{u})\,\nu_{q}(\mathbf{X}_{u})+m^{4}(\mathbf{X}_{u})\big)\Big]-\eta_{u}^{2},$ (12)
where
$\nu_{q}(\mathbf{X}_{u})=\mathbb{E}_{p}\big[w_{-u\mid u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\big]-m^{2}(\mathbf{X}_{u}).$
This expression shows that the asymptotic variance depends explicitly on $q$ through the weights $w_{u}$ and $w_{-u\mid u}$, and can be reduced by optimising them, as will be shown in the next section.
@@ 3.3 Optimal distribution: sequential construction
We aim to minimise $\sigma^{2}_{\mathrm{opt},u}(q)$ by optimising the components of $q$ in a sequential manner. The procedure involves two steps. First, for a fixed marginal $q_{u}$, we determine the optimal conditional density $q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})$ that minimises the variance contribution given $\mathbf{x}_{u}$, which is then substituted into the variance expression. In the second step, the resulting objective is minimised with respect to $q_{u}$.
To characterize the optimal distribution $q^{*}$, a principled approach consists in using the calculus of variations. Following the classical Lagrangian minimisation framework, one can formally derive the stationary condition satisfied by $q^{*}$ under the normalisation constraint. In general, verifying that this stationary point corresponds to a minimum would require checking second-order optimality conditions. However, as mentioned by Owen [34, End Notes, p. 38], the Cauchy–Schwarz inequality is sufficient to establish global minimality in this setting. For simplicity, we rely on the following key variational lemma, a standard result in importance sampling theory.
Lemma 3.3
@@
Let $\mathcal{D}_{k}$ denote the set of probability density functions on $\mathbb{R}^{k}$. Then, for any $g\in\mathcal{D}_{k}$, the functional
$J(q)=\int_{\mathbb{R}^{k}}\frac{g^{2}(\mathbf{x})}{q(\mathbf{x})}\,d\mathbf{x},\qquad q\in\mathcal{D}_{k}.$
is uniquely minimised by $g$ a.e.
Proof 1
We apply Cauchy–Schwarz inequality:
$J(g)=1=\left(\int g(\mathbf{x})\,d\mathbf{x}\right)^{2}=\left(\int\frac{g(\mathbf{x})}{q(\mathbf{x})}\,q(\mathbf{x})\,d\mathbf{x}\right)^{2}\leq\int\left(\frac{g(\mathbf{x})}{q(\mathbf{x})}\right)^{2}q(\mathbf{x})\,d\mathbf{x}=J(q),$
Equality in the Cauchy–Schwarz inequality holds if and only if the argument of the convex function is constant almost everywhere with respect to $q(\mathbf{x})\,d\mathbf{x}$. That is, there exists a constant $c>0$ such that $g(\mathbf{x})/q(\mathbf{x})=c$ a.e. Since both $g$ and $q$ integrate to one, necessarily $c=1$, and hence $q=g$ almost everywhere.
First step: optimisation of $q_{-u\mid u}$
@@
We optimise the conditional component $q_{-u\mid u}$ of $q$ for a fixed $q_{u}$. Since $\sigma^{2}_{\mathrm{opt},u}(q)$ in Eq. (12) depends on $q_{-u\mid u}$ only through the term $\mathbb{E}_{p}\!\left[w_{-u\mid u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right],$ which appears in the definition of $\nu_{q}(\mathbf{X}_{u})$, minimising this conditional expectation is sufficient. Applying Lemma 3.3 conditionally on $\mathbf{X}_{u}$ yields the optimal conditional density
$q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})=\frac{p_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})\,\phi(\mathbf{x})}{\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}=\mathbf{x}_{u}]},$ (13)
where the denominator ensures normalisation.
@@
With this choice, the variance contribution when $q_{-u\mid u}=q^{*}_{-u\mid u}$ becomes $\big(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\big)^{2}$, instead of $\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}]$ when $\mathbf{X}_{-u}\mid\mathbf{X}_{u}\sim p_{-u\mid u}$ (i.e., $w_{-u\mid u}=1$). A full derivation is provided in the Appendix.
Since the function $x\mapsto x^{2}$ is convex, Jensen’s inequality implies
$\big(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\big)^{2}\leq\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}],$
showing that $q^{*}_{-u\mid u}$ indeed reduces this part of the variance.
@@
We now fix $q_{-u\mid u}$, and seek to minimise the asymptotic variance over $q_{u}$. For any such fixed conditional, the sampling density is of the form $q=q_{u}\,q_{-u\mid u}$, and the corresponding variance reads
$\sigma_{\mathrm{opt},u}^{2}(q)=\mathbb{E}_{p}\!\left[w_{u}(\mathbf{X}_{u})\,S(\mathbf{X}_{u})\right]-\eta_{u}^{2},$
where the integrand $S(\mathbf{x}_{u})$ depends on the choice of $q_{-u\mid u}$. The optimal marginal minimising this expression is given by
$q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,\sqrt{S_{q_{-u\mid u}}(\mathbf{x}_{u})}}{\mathbb{E}_{p}\!\left[\sqrt{S_{q_{-u\mid u}}(\mathbf{X}_{u})}\right]},\qquad\text{and}\qquad\sigma^{2}_{\mathrm{opt},u}(q_{u}^{*})=\left(\mathbb{E}_{p}\!\left[\sqrt{S_{q_{-u\mid u}}(\mathbf{X}_{u})}\right]\right)^{2}-\eta_{u}^{2}.$ (14)
@@
  • Case A – Reference conditional: $q_{-u\mid u}=p_{-u\mid u}$
    $S_{A}(\mathbf{x}_{u})=4\,m^{2}(\mathbf{x}_{u})\,\nu_{p}(\mathbf{x}_{u})+m^{4}(\mathbf{x}_{u})$
  • Case B – Optimal conditional: $q_{-u\mid u}=q^{*}_{-u\mid u}$
    $S_{B}(\mathbf{x}_{u})=4\,m^{2}(\mathbf{x}_{u})\,\nu_{q^{*}}(\mathbf{x}_{u})+m^{4}(\mathbf{x}_{u})=4\,m^{2}(\mathbf{x}_{u})\big(\mathbb{E}_{p}\!\left[\phi(\mathbf{X})\mid\mathbf{X}_{u}=\mathbf{x}_{u}\right]\big)^{2}-3m^{4}(\mathbf{x}_{u}),$
Both $S_{A}(\mathbf{x}_{u})$ and $S_{B}(\mathbf{x}_{u})$ are non-negative since they are built from conditional variances $\nu_{p}(\mathbf{X}_{u})$ and $\nu_{q}(\mathbf{X}_{u})$.
@@
This satisfies the pointwise inequality $S_{B}(\mathbf{x}_{u})\leq S_{A}(\mathbf{x}_{u})$, as shown in the first step. The full chain of inequalities then reads
$\sigma^{2}_{\mathrm{opt},u}(q^{*})\leq\sigma^{2}_{\mathrm{opt},u}(q_{u}^{*}\,p_{-u\mid u})\leq\sigma^{2}_{\mathrm{opt},u}(p),$
which follows from Jensen’s inequality:
$\left(\mathbb{E}_{p}\!\left[\sqrt{S_{B}(\mathbf{X}_{u})}\right]\right)^{2}\leq\left(\mathbb{E}_{p}\!\left[\sqrt{S_{A}(\mathbf{X}_{u})}\right]\right)^{2}\leq\mathbb{E}_{p}\!\left[S_{A}(\mathbf{X}_{u})\right].$
@@
Joint optimisation in Case B therefore achieves the greatest variance reduction, demonstrating the improved efficiency obtained when both components of $q$ are optimised together.
3.4 Zero variance optimal distribution
@@
In the idealised deterministic case, where all inputs are controllable and no external noise variable $\mathbf{W}$ is present ($Y=f(\mathbf{X})$), it is theoretically possible to achieve zero variance in estimating $\eta_{u}$. This case is a direct simplification of the general framework, with the same two-step procedure but simpler expressions. We still proceed in two steps, but the expressions simplify significantly. In particular, the asymptotic variance now takes the form
$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\!\left[w_{u}(\mathbf{X}_{u})\left(4\,m^{2}(\mathbf{X}_{u})\,\mathbb{E}_{p}\!\left[w_{-u\mid u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})\,Y^{2}\,\big|\,\mathbf{X}_{u}\right]-3\,m^{4}(\mathbf{X}_{u})\right)\right]-\eta_{u}^{2}.$ (15)
Applying Lemma 3.3 sequentially yields
$q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})=\frac{p_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})\,|f(\mathbf{x})|}{\mathbb{E}_{p}\!\left[|f(\mathbf{X})|\,\big|\,\mathbf{X}_{u}=\mathbf{x}_{u}\right]},\qquad q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,\sqrt{S_{C}(\mathbf{x}_{u})}}{\mathbb{E}_{p}\!\left[\sqrt{S_{C}(\mathbf{X}_{u})}\right]},$ (16)
where
$S_{C}(\mathbf{x}_{u})=4\,m^{2}(\mathbf{x}_{u})\,\Big(\mathbb{E}_{p}\!\left[|f(\mathbf{X})|\,\big|\,\mathbf{X}_{u}=\mathbf{x}_{u}\right]\Big)^{2}-3\,m^{4}(\mathbf{x}_{u}).\\$
Hence, the optimal joint sampling distribution is $q^{*}(\mathbf{x})=q_{u}^{*}(\mathbf{x}_{u})\,q_{-u\mid u}^{*}(\mathbf{x}_{-u}\mid\mathbf{x}_{u}),$ from which we obtain
$\sigma^{2}_{\mathrm{opt},u}(q^{*})\leq\sigma^{2}_{\mathrm{opt},u}(p).$
In general, this distribution does not yield a zero variance estimator, since a direct plug-in calculation shows
$\sigma^{2}_{\mathrm{opt},u}(q^{*})=\left(\mathbb{E}_{p}\!\left[\sqrt{S_{C}(\mathbf{X}_{u})}\right]\right)^{2}-\eta_{u}^{2}>0.$
Remark. When the model $f(\mathbf{x})$ may take both positive and negative values, the theoretical zero variance density derived above is no longer a valid probability distribution. A practical alternative consists in separating the positive and negative contributions of $f$, so that each part can be sampled using its own importance distribution. This idea, sometimes referred to as positivisation in the importance sampling literature [34], relies on defining two auxiliary densities associated with the positive and negative parts of $f$.
We decompose $f$ into its positive and negative parts:
$f_{+}(\mathbf{x})=\max(f(\mathbf{x}),0),\qquad f_{-}(\mathbf{x})=\max(-f(\mathbf{x}),0),\qquad f(\mathbf{x})=f_{+}(\mathbf{x})-f_{-}(\mathbf{x}).$
Two auxiliary densities $q_{+}$ and $q_{-}$ are then defined to be positive wherever $pf_{+}$ and $pf_{-}$ are positive, respectively. Independent samples $\mathbf{X}_{i,+}\sim q_{+}$ and $\mathbf{X}_{i,-}\sim q_{-}$ are then combined with opposite signs, preserving unbiasedness while reducing variance in practice.
If $f(\mathbf{x})\geq 0$ for all $\mathbf{x}$, then $\mathbb{E}_{p}[|f(\mathbf{X})|\mid\mathbf{X}_{u}]=m(\mathbf{X}_{u})$ and $S_{C}(\mathbf{x}_{u})=m^{4}(\mathbf{x}_{u})$, so that
$q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})=\frac{p_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})\,f(\mathbf{x})}{m(\mathbf{x}_{u})},\qquad q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,m^{2}(\mathbf{x}_{u})}{\eta_{u}}.$ (17)
@@
$q^{*}(\mathbf{x})=q_{u}^{*}(\mathbf{x}_{u})\,q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})=\frac{p(\mathbf{x})\,f(\mathbf{x})\,m(\mathbf{x}_{u})}{\eta_{u}},\qquad\text{and}\qquad\sigma^{2}_{\mathrm{opt},u}(q^{*})=0.$ (18)
In theory, the density $q^{*}(\mathbf{x})$ yields a zero variance estimator, but this ideal case cannot be achieved in practice since the optimal distribution depends on the unknown target quantity $\eta_{u}$, which brings us back to the classical importance sampling problem of approximating it. Adaptive procedures can be employed, including nonparametric methods [49] and parametric ones such as the cross-entropy method [38, 14].